### v1

# DRESS: A Continuous Framework for Structural Graph Refinement

Abstract
The Weisfeiler-Lehman (WL) hierarchy is a cornerstone framework for graph isomorphism testing and structural analysis. However, scaling beyond 1-WL to 3-WL and higher requires tensor-based operations that scale as $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$, making them computationally prohibitive for large graphs. In this paper, we start from the Original-DRESS equation (Castrillo, León, and Gómez, 2018)—a parameter-free, continuous dynamical system on edges—and show that it distinguishes the prism graph from $K_{3,3}$, a pair that 1-WL provably cannot separate. We then generalize it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions, and further to Generalized-DRESS, an abstract template parameterized by the choice of neighborhood operator, aggregation function and norm. Finally, we introduce $\Delta$-DRESS, which runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the Kelly–Ulam reconstruction conjecture. Both Motif-DRESS and $\Delta$-DRESS empirically distinguish Strongly Regular Graphs (SRGs)—such as the Rook and Shrikhande graphs—that confound 3-WL. Our results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL on well-known benchmark graphs, without the prohibitive $\mathcal{O}(n^{4})$ computational cost.
1 Introduction
The graph isomorphism problem and structural similarity measurement rely heavily on the Weisfeiler-Lehman (WL) hierarchy [7]. While the standard 1-WL test—shown to characterize the expressiveness of message-passing neural networks [8, 6]—is highly efficient, it fails to distinguish many non-isomorphic graphs. Higher-order $k$-WL tests [2] offer greater expressiveness, but their $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$ complexity renders them impractical for large-scale network analysis.
The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation. As we show in Section 3, Original-DRESS distinguishes graphs that 1-WL cannot. However, it still fails to distinguish highly symmetric structures such as Strongly Regular Graphs (SRGs).
In this work, we present a hierarchy of continuous structural refinement algorithms, building from the concrete to the general:
  1. 1.
    Original-DRESS (Section 2): The foundational equation, a parameter-free dynamical system on edges.
  2. 2.
    Original-DRESS vs. 1-WL (Section 3): We show that Original-DRESS distinguishes graphs that 1-WL cannot, including the prism graph from $K_{3,3}$.
  3. 3.
    Motif-DRESS (Section 4): A generalization that replaces triangle neighborhoods with arbitrary structural motifs. Original-DRESS is recovered as a special case with $M=K_{3}$.
  4. 4.
    Generalized-DRESS (Section 5): The most abstract template, opening the aggregation function and norm as additional free parameters.
  5. 5.
    $\Delta$-DRESS (Section 6): Runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the reconstruction conjecture and empirically distinguishing SRGs that confound 3-WL.
We show that Original-DRESS distinguishes graphs that 1-WL cannot, and demonstrate empirically that Motif-DRESS and $\Delta$-DRESS distinguish graphs beyond the reach of 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs.
Isomorphism Test.
Throughout this paper, the graph fingerprint is the sorted vector of converged edge values $\operatorname{sort}(d^{*})$. Two graphs are declared non-isomorphic if and only if their fingerprints differ element-wise beyond a numerical tolerance $\epsilon$. This applies uniformly to all DRESS variants.
All experiments use a convergence tolerance of $\epsilon=10^{-6}$ and a maximum of 20 iterations; in every case convergence was reached before the iteration limit.
Related Work.
Higher-order GNN architectures that exceed 1-WL expressiveness include $k$-GNN [6], GIN [8], PPGN [5], and Ring-GNN [4]. Subgraph-based approaches such as GNN-AK+ [9] and ESAN [1] achieve higher expressiveness by running message-passing networks on node-deleted or node-marked subgraphs; $\Delta$-DRESS (Section 6) shares the node-deletion strategy but replaces learned message passing with deterministic fixed-point iteration. All of the above are supervised methods that require labeled training data and learnable parameters. In contrast, DRESS is a deterministic, unsupervised framework that produces canonical fingerprints via fixed-point iteration, without any learning or parameter tuning.
2 Original-DRESS
The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation:
$d_{uv}^{(t+1)}=\frac{\sum_{x\in N[u]\cap N[v]}\bigl(d_{ux}^{(t)}+d_{xv}^{(t)}\bigr)}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (1)
where $N[u]=N(u)\cup\{u\}$ denotes the closed neighborhood of $u$, and $\|u\|=\sqrt{\sum_{x\in N[u]}d_{ux}^{(t)}}$ is the node norm. The self-similarity $d_{uu}=2$ is invariant under iteration. This equation converges to a unique fixed point, providing a continuous structural fingerprint for the graph.
Self-loops. Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$). The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbours would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
Initialization. The equation is scale-invariant (degree-0 homogeneous), so the fixed point is independent of the initial scale. Any uniform initialisation $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$. In practice, $c=1$ is used.
However, because Original-DRESS aggregates strictly over triangles (common neighbours), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
We now demonstrate a key expressiveness result: Original-DRESS—the simplest member of the DRESS family—already distinguishes graphs that 1-WL (color refinement) cannot.
Theorem 1 (DRESS distinguishes beyond 1-WL).
There exist graph pairs that 1-WL cannot distinguish but Original-DRESS can. In particular, DRESS distinguishes the prism graph ($C_{3}\square K_{2}$) from the complete bipartite graph $K_{3,3}$, a pair that 1-WL provably cannot separate.
Proof.
Both graphs are 3-regular on 6 vertices with 9 edges. Since all vertices have the same degree, 1-WL assigns a uniform color to every vertex in both graphs after every iteration, and therefore cannot distinguish them [2].
DRESS operates on edges and aggregates over common neighbours (triangles). We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
$K_{3,3}$: No two adjacent vertices share a common neighbour ($K_{3,3}$ is triangle-free). Hence every edge has the same neighborhood structure, and by symmetry the unique fixed point assigns the same value to all 9 edges. Numerically, $d^{*}_{K_{3,3}}\approx 1.155$ for all edges.
Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbour) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbours besides self-loops). Suppose for contradiction that all 9 edges converge to the same value $d^{*}$. Then:
  • Triangle edges: $d^{*}=\dfrac{4+4d^{*}}{2+3d^{*}}$, which gives $d^{*}=\frac{1+\sqrt{13}}{3}\approx 1.535$.
  • Matching edges: $d^{*}=\dfrac{4+2d^{*}}{2+3d^{*}}$, which gives $d^{*}=\frac{2}{\sqrt{3}}\approx 1.155$.
These are contradictory, so the prism cannot have a uniform fixed point. The converged unique edge values are:
Graph Unique values Multiplicities
Prism $\{0.922,\;1.709\}$ 3 matching edges, 6 triangle edges
$K_{3,3}$ $\{2/\sqrt{3}\approx 1.155\}$ all 9 edges identical
Since 1-WL cannot distinguish them and DRESS can, DRESS distinguishes beyond 1-WL on this instance. ∎
Remark.
This result holds for Original-DRESS alone—the simplest member of the DRESS family, using only triangle neighborhoods and $\mathcal{O}(E)$ computation per iteration. The generalizations that follow (Motif-DRESS, $\Delta$-DRESS) extend this advantage further to graphs that resist even 3-WL.
4 Motif-DRESS
Original-DRESS is limited to triangle neighborhoods. What if we use other structural motifs? Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional weight function $w(e)$:
$d_{uv}^{(t+1)}=\frac{\displaystyle\sum_{e^{\prime}\in\mathcal{N}_{M}(u,v)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (2)
where $\|u\|^{(t)}=\sqrt{\sum_{e^{\prime}\in\mathcal{N}_{M}(u,u)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}$ is the weighted node norm at time $t$.
The weights $w(e)$ act as a multiplicative factor, controlling how much structural information flows along each edge. Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
Different choices of motif $M$ yield different neighborhood operators:
  • Triangle ($M=K_{3}$): $\mathcal{N}(u,v)=\{(u,x),(x,v)\mid x\in N[u]\cap N[v]\}$—this recovers Original-DRESS.
  • 4-cycle ($M=C_{4}$): Aggregates over edges participating in 4-cycles with $(u,v)$.
  • $K_{4}$ clique: Aggregates over edges forming a $K_{4}$ with $(u,v)$.
4.1 Properties
Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists. For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot\Delta)$. For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
Invariant: $d_{uu}=2$. The self-similarity is invariant under iteration: since $\mathcal{N}_{M}(u,u)$ includes all edges incident to $u$ symmetrically, and $d_{ux}=d_{xu}$, substituting $v=u$ yields $d_{uu}^{(t+1)}=2\|u\|^{2}/\|u\|^{2}=2$ for all $t$. Thus $d_{uu}=2$ is a fixed property of the equation, not a free parameter.
4.2 Expressiveness
All experiments below use the $K_{4}$ clique motif. The specific SRG pairs tested below are known to be indistinguishable by 3-WL; each successful distinction therefore demonstrates that Motif-DRESS empirically exceeds 3-WL on these instances.
  • Rook vs. Shrikhande: Successfully distinguishes this pair of SRGs with parameters $(16,6,2,2)$. The Rook graph ($K_{4}\square K_{4}$) contains $K_{4}$ cliques while the Shrikhande graph does not, so the $K_{4}$-neighborhood sizes differ per edge.
  • Chang Graphs: Distinguishes all 6 pairwise comparisons among the four SRGs with parameters $(28,12,6,4)$: T(8), Chang-1, Chang-2, and Chang-3.
4.3 Proof Sketch of Convergence
Theorem 2 (Convergence of Motif-DRESS).
Let $F:\mathbb{R}_{+}^{|E|}\to\mathbb{R}_{+}^{|E|}$ be the Motif-DRESS update operator. If the aggregation function and the node norm satisfy the degree-0 homogeneity condition ($p=2q$), then $F$ converges to a unique fixed point $d^{*}\in[0,2]^{|E|}$ for any non-negative initial state $d^{(0)}>0$.
Proof sketch.
The argument proceeds in three steps:
  1. 1.
    Scale Invariance (Degree-0 Homogeneity): The aggregation function is positively homogeneous of degree $p$, and the norm of degree $q$. Since $p=2q$, the ratio $F(d)$ is homogeneous of degree 0: for any $\lambda>0$, $F(\lambda d)=F(d)$. Hence the iteration is self-regularizing and cannot diverge or collapse to zero.
  2. 2.
    Boundedness: For any valid symmetric neighborhood operator $\mathcal{N}_{M}(e)$, the Cauchy–Schwarz inequality (or Hölder’s inequality for Minkowski-$r$ variants) bounds the numerator by the denominator. Self-loop inclusion ensures the denominator is strictly positive. Thus $F(d)_{uv}\in[0,2]$ for all $d>0$.
  3. 3.
    Contraction on the Hilbert Projective Metric: Since $F$ is a positive, degree-0 homogeneous map on the cone $\mathbb{R}_{>0}^{|E|}$, Birkhoff’s contraction theorem guarantees that $F$ is a strict contraction under the Hilbert projective metric $d_{H}(x,y)=\log\!\bigl(\max_{e}x_{e}/y_{e}\cdot\max_{e}y_{e}/x_{e}\bigr)$, provided $F$ maps a bounded part of the cone into a strictly smaller part—which follows from the Cauchy–Schwarz bound above. By the Banach fixed-point theorem on the complete metric space $(\mathbb{R}_{>0}^{|E|}/\sim,\,d_{H})$, the iteration converges to a unique ray, and the boundedness step pins the representative to $d^{*}\in[0,2]^{|E|}$.
A complete formal verification of the contraction constant is deferred to future work; all empirical tests confirm convergence within 20 iterations. ∎
5 Generalized-DRESS
Motif-DRESS fixes the aggregation to summation and the norm to the geometric mean. Generalized-DRESS is the most abstract template, allowing any choice of these components. For each edge $(u,v)$ and a symmetric neighborhood operator $\mathcal{N}(u,v)$:
$d_{uv}^{(t+1)}=\frac{f\bigl(\{d_{e^{\prime}}^{(t)}\}_{e^{\prime}\in\mathcal{N}(u,v)}\bigr)}{g^{(t)}(u)\cdot g^{(t)}(v)}$ (3)
where:
  • $\mathcal{N}(u,v)$ is a symmetric neighborhood operator—the set of edges aggregated for edge $(u,v)$,
  • $f$ is the aggregation function over edges in $\mathcal{N}(u,v)$,
  • $g(u)$ is the node norm.
Original-DRESS and Motif-DRESS are both special cases: Original-DRESS fixes $\mathcal{N}$ to triangles, $f=\text{sum}$, $g=\text{geometric norm}$; Motif-DRESS generalizes $\mathcal{N}$ to arbitrary motifs while keeping the same $f$ and $g$. Generalized-DRESS opens all three parameters, enabling variants such as Cosine-DRESS (cosine similarity aggregation) or Minkowski-$r$ norms.
6 $\Delta$-DRESS
$\Delta$-DRESS breaks symmetry by running DRESS on each node-deleted subgraph $G\setminus\{v\}$ for every $v\in V$. The graph fingerprint is the sorted multiset of $n$ converged edge-value vectors (one per deletion), compared by flattening and sorting without any summarization.
Deleting a node from a regular graph produces an irregular subgraph where DRESS can now distinguish structure that was hidden by the uniform regularity.
6.1 Connection to the Reconstruction Conjecture
The multiset $\{\!\{\text{DRESS}(G\setminus\{v\}):v\in V\}\!\}$ is directly analogous to the deck in the Kelly–Ulam reconstruction conjecture, which posits that graphs with $n\geq 3$ are determined (up to isomorphism) by their multiset of node-deleted subgraphs. $\Delta$-DRESS computes a continuous relaxation of this deck.
6.2 Properties
Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot\Delta)$, where $I$ is the number of iterations per deletion. The computation for each deleted subgraph is entirely independent, making $\Delta$-DRESS embarrassingly parallel.
6.3 Expressiveness
$\Delta$-DRESS empirically distinguishes the following non-isomorphic pairs:
  • Rook vs. Shrikhande: Successfully distinguished (SRG(16,6,2,2)(16,6,2,2); confounds 3-WL).
  • Chang Graphs: Distinguished all 6 pairs among T(8) and the three Chang graphs (SRG(28,12,6,4)(28,12,6,4); confound 3-WL).
  • $2\times C_{4}$ vs. $C_{8}$: Successfully distinguished (both 2-regular on 8 nodes).
  • Petersen vs. Pentagonal Prism: Successfully distinguished (both 3-regular on 10 nodes).
7 Family Structure
The DRESS variants introduced in this paper form a nested hierarchy with an orthogonal composition operator:
$\text{Generalized-DRESS}\supset\text{Motif-DRESS}\supset\text{Original-DRESS}$
$\Delta(\cdot)$ is an orthogonal wrapper applicable to any of the above: given any DRESS variant $\mathcal{F}$,
$\Delta\text{-DRESS}(\mathcal{F})=\{\!\{\mathcal{F}(G\setminus\{v\}):v\in V\}\!\}$
The node-deletion strategy is independent of the choice of $\mathcal{N}$, $f$, and $g$. All experiments in Section 6 use $\mathcal{F}=\text{Original-DRESS}$, but $\Delta$-Motif-DRESS or $\Delta$-Cosine-DRESS are equally valid and may offer complementary expressiveness. The deletion strategy generalizes naturally to $\Delta^{k}$-DRESS, which runs DRESS on all $\binom{n}{k}$ node-deleted subgraphs; we leave the exploration of this hierarchy to future work.
8 Conclusion
We presented the DRESS family, a hierarchy of continuous structural refinement algorithms. Starting from Original-DRESS—a single equation proposed in 2018—we showed it distinguishes graphs that 1-WL cannot (Theorem 1). We then generalized it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions. Generalized-DRESS abstracts further, opening the aggregation function and norm as free parameters. $\Delta$-DRESS extends the framework by running DRESS on each node-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture and empirically distinguishing Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs. These results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL, without the prohibitive computational cost of higher-order WL tests. An open-source implementation with bindings for C, C++, Python, Rust, Go, Julia, R, MATLAB, and WebAssembly is publicly available at https://github.com/velicast/dress-graph.

### v2

@@
-DRESS: A Continuous Framework for Structural Graph Refinement
+arXiv:2602.20833v2[cs.DS] 26 Feb 2026
 Abstract
@@
 Higher-order $k$-WL tests [2] offer greater expressiveness, but their $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$ complexity renders them impractical for large-scale network analysis.
-The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation.
+The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbor aggregation.
 As we show in Section 3, Original-DRESS distinguishes graphs that 1-WL cannot.
@@
 This applies uniformly to all DRESS variants.
+An equivalent representation is the histogram $h(d^{*})$: since all DRESS values lie in $[0,2]$ and the convergence tolerance is $\epsilon=10^{-6}$, each value maps to one of $\lceil 2/\epsilon\rceil=2\times 10^{6}$ integer bins, yielding a fixed-size bin-count vector that uniquely identifies the same multiset.
+In general, the number of bins is determined by the convergence tolerance: a tolerance of $\epsilon$ over the range $[0,2]$ gives $\lceil 2/\epsilon\rceil$ bins.
+The histogram is particularly useful when fingerprints from multiple subgraphs must be pooled (see $\Delta$-DRESS, Section 6).
 All experiments use a convergence tolerance of $\epsilon=10^{-6}$ and a maximum of 20 iterations; in every case convergence was reached before the iteration limit.
@@
 2 Original-DRESS
-The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation:
+The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbor aggregation.
 $d_{uv}^{(t+1)}=\frac{\sum_{x\in N[u]\cap N[v]}\bigl(d_{ux}^{(t)}+d_{xv}^{(t)}\bigr)}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (1)
@@
 Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$).
-The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbours would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
+The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbors would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
 Initialization.
 The equation is scale-invariant (degree-0 homogeneous), so the fixed point is independent of the initial scale.
-Any uniform initialisation $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$.
+Any uniform initialization $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$.
 In practice, $c=1$ is used.
-However, because Original-DRESS aggregates strictly over triangles (common neighbours), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
+However, because Original-DRESS aggregates strictly over triangles (common neighbors), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
 3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
@@
 Since all vertices have the same degree, 1-WL assigns a uniform color to every vertex in both graphs after every iteration, and therefore cannot distinguish them [2].
-DRESS operates on edges and aggregates over common neighbours (triangles).
-We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
-$K_{3,3}$: No two adjacent vertices share a common neighbour ($K_{3,3}$ is triangle-free).
+DRESS operates on edges and aggregates over common neighbors (triangles). We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
+$K_{3,3}$: No two adjacent vertices share a common neighbor ($K_{3,3}$ is triangle-free).
 Hence every edge has the same neighborhood structure, and by symmetry the unique fixed point assigns the same value to all 9 edges.
 Numerically, $d^{*}_{K_{3,3}}\approx 1.155$ for all edges.
-Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbour) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbours besides self-loops).
+Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbor) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbors besides self-loops).
 Suppose for contradiction that all 9 edges converge to the same value $d^{*}$.
@@
 What if we use other structural motifs?
-Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional weight function $w(e)$:
+Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional symmetric weight function $w:E\to\mathbb{R}_{>0}$:
 $d_{uv}^{(t+1)}=\frac{\displaystyle\sum_{e^{\prime}\in\mathcal{N}_{M}(u,v)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (2)
 where $\|u\|^{(t)}=\sqrt{\sum_{e^{\prime}\in\mathcal{N}_{M}(u,u)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}$ is the weighted node norm at time $t$.
-The weights $w(e)$ act as a multiplicative factor, controlling how much structural information flows along each edge.
+The symmetric weight function $w(e)$ acts as a multiplicative factor, controlling how much structural information flows along each edge.
 Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
@@
 Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists.
-For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot\Delta)$.
+For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot T)$, where $T$ is the number of triangles.
 For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
@@
 By the Banach fixed-point theorem on the complete metric space $(\mathbb{R}_{>0}^{|E|}/\sim,\,d_{H})$, the iteration converges to a unique ray, and the boundedness step pins the representative to $d^{*}\in[0,2]^{|E|}$.
-A complete formal verification of the contraction constant is deferred to future work; all empirical tests confirm convergence within 20 iterations. ∎
+A complete formal verification of the contraction constant is deferred to future work; all empirical tests in this paper confirm convergence within 20 iterations. ∎
 5 Generalized-DRESS
@@
 6.2 Properties
-Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot\Delta)$, where $I$ is the number of iterations per deletion.
+Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot T)$, where $T$ is the number of triangles and $I$ is the number of iterations per deletion.
 The computation for each deleted subgraph is entirely independent, making $\Delta$-DRESS embarrassingly parallel.

### v3

@@
-arXiv:2602.20833v2[cs.DS] 26 Feb 2026
+arXiv:2602.20833v3[cs.DS] 02 Mar 2026
 Abstract
@@
 However, scaling beyond 1-WL to 3-WL and higher requires tensor-based operations that scale as $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$, making them computationally prohibitive for large graphs.
-In this paper, we start from the Original-DRESS equation (Castrillo, León, and Gómez, 2018)—a parameter-free, continuous dynamical system on edges—and show that it distinguishes the prism graph from $K_{3,3}$, a pair that 1-WL provably cannot separate.
+In this paper, we start from the Original-DRESS equation (Castrillo, León, and Gómez, 2018)–a parameter-free, continuous dynamical system on edges–and show that it distinguishes the prism graph from $K_{3,3}$, a pair that 1-WL provably cannot separate.
 We then generalize it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions, and further to Generalized-DRESS, an abstract template parameterized by the choice of neighborhood operator, aggregation function and norm.
-Finally, we introduce $\Delta$-DRESS, which runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the Kelly–Ulam reconstruction conjecture.
-Both Motif-DRESS and $\Delta$-DRESS empirically distinguish Strongly Regular Graphs (SRGs)—such as the Rook and Shrikhande graphs—that confound 3-WL.
+Finally, we introduce $\Delta$-DRESS, which runs DRESS on each vertex-deleted subgraph $G\setminus\{v\}$, connecting the framework to the Kelly–Ulam reconstruction conjecture.
+$\Delta$-DRESS empirically distinguishes Strongly Regular Graphs (SRGs)–such as the Rook and Shrikhande graphs–that confound 3-WL.
 Our results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL on well-known benchmark graphs, without the prohibitive $\mathcal{O}(n^{4})$ computational cost.
@@
 The graph isomorphism problem and structural similarity measurement rely heavily on the Weisfeiler-Lehman (WL) hierarchy [8].
-While the standard 1-WL test—shown to characterize the expressiveness of message-passing neural networks [8, 6]—is highly efficient, it fails to distinguish many non-isomorphic graphs.
+While the standard 1-WL test–shown to characterize the expressiveness of message-passing neural networks [9, 7]–is highly efficient, it fails to distinguish many non-isomorphic graphs.
 Higher-order $k$-WL tests [2] offer greater expressiveness, but their $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$ complexity renders them impractical for large-scale network analysis.
@@
   5. 5.
-    $\Delta$-DRESS (Section 6): Runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the reconstruction conjecture and empirically distinguishing SRGs that confound 3-WL.
-We show that Original-DRESS distinguishes graphs that 1-WL cannot, and demonstrate empirically that Motif-DRESS and $\Delta$-DRESS distinguish graphs beyond the reach of 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs.
+    $\Delta$-DRESS (Section 6): Runs DRESS on each vertex-deleted subgraph $G\setminus\{v\}$, connecting the framework to the reconstruction conjecture.
 Isomorphism Test.
@@
 This applies uniformly to all DRESS variants.
-An equivalent representation is the histogram $h(d^{*})$: since all DRESS values lie in $[0,2]$ and the convergence tolerance is $\epsilon=10^{-6}$, each value maps to one of $\lceil 2/\epsilon\rceil=2\times 10^{6}$ integer bins, yielding a fixed-size bin-count vector that uniquely identifies the same multiset.
+An equivalent representation is the histogram $h(d^{*})$: in the unweighted case all DRESS values lie in $[0,2]$ (see the convergence proof in Section 4.3) and the convergence tolerance is $\epsilon=10^{-6}$, so each value maps to one of $\lceil 2/\epsilon\rceil=2\times 10^{6}$ integer bins, yielding a fixed-size bin-count vector that uniquely identifies the same multiset.
 In general, the number of bins is determined by the convergence tolerance: a tolerance of $\epsilon$ over the range $[0,2]$ gives $\lceil 2/\epsilon\rceil$ bins.
+For weighted graphs, where values may exceed $2$, the bin range is extended accordingly.
 The histogram is particularly useful when fingerprints from multiple subgraphs must be pooled (see $\Delta$-DRESS, Section 6).
@@
 Higher-order GNN architectures that exceed 1-WL expressiveness include $k$-GNN [7], GIN [9], PPGN [6], and Ring-GNN [5].
-Subgraph-based approaches such as GNN-AK+ [9] and ESAN [1] achieve higher expressiveness by running message-passing networks on node-deleted or node-marked subgraphs; $\Delta$-DRESS (Section 6) shares the node-deletion strategy but replaces learned message passing with deterministic fixed-point iteration.
+Subgraph-based approaches such as GNN-AK+ [10] and ESAN [1] achieve higher expressiveness by running message-passing networks on vertex-deleted or vertex-marked subgraphs; $\Delta$-DRESS (Section 6) shares the vertex-deletion strategy but replaces learned message passing with deterministic fixed-point iteration.
 All of the above are supervised methods that require labeled training data and learnable parameters.
@@
 $d_{uv}^{(t+1)}=\frac{\sum_{x\in N[u]\cap N[v]}\bigl(d_{ux}^{(t)}+d_{xv}^{(t)}\bigr)}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (1)
-where $N[u]=N(u)\cup\{u\}$ denotes the closed neighborhood of $u$, and $\|u\|=\sqrt{\sum_{x\in N[u]}d_{ux}^{(t)}}$ is the node norm.
+where $N[u]=N(u)\cup\{u\}$ denotes the closed neighborhood of $u$, and $\|u\|=\sqrt{\sum_{x\in N[u]}d_{ux}^{(t)}}$ is the vertex norm.
 The self-similarity $d_{uu}=2$ is invariant under iteration.
@@
 Self-loops.
-Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$).
-The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbors would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
+Self-loops are added to every vertex before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$).
+The self-loop edge $(u,u)$ participates in both the aggregation and the vertex norm; without it, an isolated edge with no common neighbors would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
 Initialization.
@@
 3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
-We now demonstrate a key expressiveness result: Original-DRESS—the simplest member of the DRESS family—already distinguishes graphs that 1-WL (color refinement) cannot.
+We now demonstrate a key expressiveness result: Original-DRESS–the simplest member of the DRESS family–already distinguishes graphs that 1-WL (color refinement) cannot.
 Theorem 1 (DRESS distinguishes beyond 1-WL).
@@
 Remark.
-This result holds for Original-DRESS alone—the simplest member of the DRESS family, using only triangle neighborhoods and $\mathcal{O}(E)$ computation per iteration.
+This result holds for Original-DRESS alone–the simplest member of the DRESS family, using only triangle neighborhoods and $\mathcal{O}(E)$ computation per iteration.
 The generalizations that follow (Motif-DRESS, $\Delta$-DRESS) extend this advantage further to graphs that resist even 3-WL.
@@
 What if we use other structural motifs?
-Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional symmetric weight function $w:E\to\mathbb{R}_{>0}$:
-$d_{uv}^{(t+1)}=\frac{\displaystyle\sum_{e^{\prime}\in\mathcal{N}_{M}(u,v)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (2)
-where $\|u\|^{(t)}=\sqrt{\sum_{e^{\prime}\in\mathcal{N}_{M}(u,u)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}$ is the weighted node norm at time $t$.
+Motif-DRESS generalizes Original-DRESS by replacing the common-neighbor aggregation with a motif-defined symmetric vertex neighborhood $\mathcal{N}_{M}(u,v)=\mathcal{N}_{M}(v,u)$: the set of endpoints of edges in instances of an undirected motif $M$ containing edge $(u,v)$ that are adjacent to $u$ or $v$, always including $u$ and $v$ themselves, with an optional symmetric weight function $w:E\to\mathbb{R}_{>0}$ (with $w_{e}=1$ for unweighted graphs).
+$d_{uv}^{(t+1)}=\frac{\displaystyle\sum_{x\in\mathcal{N}_{M}(u,v)}\bigl(w_{ux}\cdot d_{ux}^{(t)}+w_{xv}\cdot d_{xv}^{(t)}\bigr)}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (2)
+where $\|u\|^{(t)}=\sqrt{\sum_{x\in N[u]}w_{ux}\cdot d_{ux}^{(t)}}$, and $w_{ab}\cdot d_{ab}=0$ whenever $(a,b)\notin E$.
 The symmetric weight function $w(e)$ acts as a multiplicative factor, controlling how much structural information flows along each edge.
 Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
-Different choices of motif $M$ yield different neighborhood operators:
-  • Triangle ($M=K_{3}$): $\mathcal{N}(u,v)=\{(u,x),(x,v)\mid x\in N[u]\cap N[v]\}$—this recovers Original-DRESS.
-  • 4-cycle ($M=C_{4}$): Aggregates over edges participating in 4-cycles with $(u,v)$.
-  • $K_{4}$ clique: Aggregates over edges forming a $K_{4}$ with $(u,v)$.
+Different choices of motif $M$ yield different neighborhoods:
+  • Triangle ($M=K_{3}$): $\mathcal{N}_{K_{3}}(u,v)=N[u]\cap N[v]$, the closed common neighborhood.
+This recovers Original-DRESS exactly: $\sum_{x\in N[u]\cap N[v]}\bigl(w_{ux}\,d_{ux}+w_{xv}\,d_{xv}\bigr)$.
+  • $K_{4}$ clique: For each pair $x,y$ with $\{u,v,x,y\}$ forming a $K_{4}$, $\mathcal{N}_{K_{4}}(u,v)$ contains $u$, $v$, $x$, and $y$. Each vertex contributes $w_{ux}\,d_{ux}+w_{xv}\,d_{xv}$.
+  • 4-cycle ($M=C_{4}$): For each 4-cycle $u$-$x$-$y$-$v$, $\mathcal{N}_{C_{4}}(u,v)$ contains $u$, $v$, $x$, and $y$.
+Vertex $x$ contributes $w_{ux}\,d_{ux}$; vertex $y$ contributes $w_{yv}\,d_{yv}$ (cross-terms are zero since $(x,v),(u,y)\notin E$ in the $C_{4}$).
 4.1 Properties
-Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists.
+Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif neighborhoods.
 For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot T)$, where $T$ is the number of triangles.
-For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
+For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif neighborhoods can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
 Invariant: $d_{uu}=2$.
-The self-similarity is invariant under iteration: since $\mathcal{N}_{M}(u,u)$ includes all edges incident to $u$ symmetrically, and $d_{ux}=d_{xu}$, substituting $v=u$ yields $d_{uu}^{(t+1)}=2\|u\|^{2}/\|u\|^{2}=2$ for all $t$.
-Thus $d_{uu}=2$ is a fixed property of the equation, not a free parameter.
+The self-similarity $d_{uu}=2$ is a constant maintained throughout iteration.
+Since the iteration only updates $d_{uv}$ for $u\neq v$, and the norm $\|u\|=\sqrt{\sum_{x\in N[u]}w_{ux}\cdot d_{ux}}$ always includes $d_{uu}=2$, this is a fixed property of the equation, not a free parameter.
 4.2 Expressiveness
@@
 The Rook graph ($K_{4}\square K_{4}$) contains $K_{4}$ cliques while the Shrikhande graph does not, so the $K_{4}$-neighborhood sizes differ per edge.
-  • Chang Graphs: Distinguishes all 6 pairwise comparisons among the four SRGs with parameters $(28,12,6,4)$: T(8), Chang-1, Chang-2, and Chang-3.
+  • Chang Graphs: Distinguishes 3 of the 6 pairwise comparisons among the four SRGs with parameters $(28,12,6,4)$: T(8) vs each of Chang-1, Chang-2, and Chang-3.
+The three Chang graphs are pairwise indistinguishable by Motif-$K_{4}$ (all three have identical $K_{4}$-neighborhood structure per edge).
 4.3 Proof Sketch of Convergence
@@
 Let $F:\mathbb{R}_{+}^{|E|}\to\mathbb{R}_{+}^{|E|}$ be the Motif-DRESS update operator.
-If the aggregation function and the node norm satisfy the degree-0 homogeneity condition ($p=2q$), then $F$ converges to a unique fixed point $d^{*}\in[0,2]^{|E|}$ for any non-negative initial state $d^{(0)}>0$.
+If the aggregation function and the vertex norm satisfy the degree-0 homogeneity condition ($p=2q$), then $F$ converges to a unique fixed point $d^{*}$ for any non-negative initial state $d^{(0)}>0$.
+In the unweighted case $d^{*}\in[0,2]^{|E|}$; with non-uniform edge weights, values may exceed $2$.
 Proof sketch.
@@
   2. 2.
-    Boundedness: For any valid symmetric neighborhood operator $\mathcal{N}_{M}(e)$, the Cauchy–Schwarz inequality (or Hölder’s inequality for Minkowski-$r$ variants) bounds the numerator by the denominator.
-Self-loop inclusion ensures the denominator is strictly positive.
-Thus $F(d)_{uv}\in[0,2]$ for all $d>0$.
+    Boundedness: The motif neighborhood $\mathcal{N}_{M}(u,v)$ is a subset of each vertex’s full neighborhood, so the numerator decomposes as $\sum_{\mathcal{N}_{M}}\bar{w}_{ux}d_{ux}+\sum_{\mathcal{N}_{M}}\bar{w}_{vx}d_{vx}\leq\|u\|^{2}+\|v\|^{2}$.
+Self-loop inclusion ensures $\|u\|,\|v\|>0$, giving the per-step bound $F(d)_{uv}\leq\|u\|/\|v\|+\|v\|/\|u\|$, which equals $2$ iff $\|u\|=\|v\|$ (AM–GM).
+In the unweighted case the fixed-point contraction forces all norms to equalize, yielding $d^{*}\in[0,2]$.
+With non-uniform edge weights, node norms may differ and converged values can exceed $2$.
   3. 3.
-    Contraction on the Hilbert Projective Metric: Since $F$ is a positive, degree-0 homogeneous map on the cone $\mathbb{R}_{>0}^{|E|}$, Birkhoff’s contraction theorem guarantees that $F$ is a strict contraction under the Hilbert projective metric $d_{H}(x,y)=\log\!\bigl(\max_{e}x_{e}/y_{e}\cdot\max_{e}y_{e}/x_{e}\bigr)$, provided $F$ maps a bounded part of the cone into a strictly smaller part—which follows from the Cauchy–Schwarz bound above.
-By the Banach fixed-point theorem on the complete metric space $(\mathbb{R}_{>0}^{|E|}/\sim,\,d_{H})$, the iteration converges to a unique ray, and the boundedness step pins the representative to $d^{*}\in[0,2]^{|E|}$.
+    Contraction on the Hilbert Projective Metric: Since $F$ is a positive, degree-0 homogeneous map on the cone $\mathbb{R}_{>0}^{|E|}$, Birkhoff’s contraction theorem guarantees that $F$ is a strict contraction under the Hilbert projective metric $d_{H}(x,y)=\log\!\bigl(\max_{e}x_{e}/y_{e}\cdot\max_{e}y_{e}/x_{e}\bigr)$, provided $F$ maps a bounded part of the cone into a strictly smaller part–which follows from the finite per-step bound above.
+By the Banach fixed-point theorem on the complete metric space $(\mathbb{R}_{>0}^{|E|}/\sim,\,d_{H})$, the iteration converges to a unique ray, and the boundedness step pins the representative to a finite vector $d^{*}$ (in the unweighted case, $d^{*}\in[0,2]^{|E|}$).
 A complete formal verification of the contraction constant is deferred to future work; all empirical tests in this paper confirm convergence within 20 iterations. ∎
@@
 Motif-DRESS fixes the aggregation to summation and the norm to the geometric mean.
-Generalized-DRESS is the most abstract template, allowing any choice of these components.
-For each edge $(u,v)$ and a symmetric neighborhood operator $\mathcal{N}(u,v)$:
+Generalized-DRESS is the most abstract template, allowing any choice of these components as long as the resulting update rule preserves the convergence guarantees (degree-0 homogeneity, boundedness, and contraction; see Section 4.3).
+For each edge $(u,v)$:
 $d_{uv}^{(t+1)}=\frac{f\bigl(\{d_{e^{\prime}}^{(t)}\}_{e^{\prime}\in\mathcal{N}(u,v)}\bigr)}{g^{(t)}(u)\cdot g^{(t)}(v)}$ (3)
 where:
-  • $\mathcal{N}(u,v)$ is a symmetric neighborhood operator—the set of edges aggregated for edge $(u,v)$,
+  • $\mathcal{N}(u,v)$ is a symmetric neighborhood operator–the set of edges aggregated for edge $(u,v)$,
   • $f$ is the aggregation function over edges in $\mathcal{N}(u,v)$,
-  • $g(u)$ is the node norm.
+  • $g(u)$ is the vertex norm.
 Original-DRESS and Motif-DRESS are both special cases: Original-DRESS fixes $\mathcal{N}$ to triangles, $f=\text{sum}$, $g=\text{geometric norm}$; Motif-DRESS generalizes $\mathcal{N}$ to arbitrary motifs while keeping the same $f$ and $g$.
 Generalized-DRESS opens all three parameters, enabling variants such as Cosine-DRESS (cosine similarity aggregation) or Minkowski-$r$ norms.
-6 $\Delta$-DRESS
-$\Delta$-DRESS breaks symmetry by running DRESS on each node-deleted subgraph $G\setminus\{v\}$ for every $v\in V$.
-The graph fingerprint is the sorted multiset of $n$ converged edge-value vectors (one per deletion), compared by flattening and sorting without any summarization.
-Deleting a node from a regular graph produces an irregular subgraph where DRESS can now distinguish structure that was hidden by the uniform regularity.
+6 $\Delta$-DRESS: Deletion
+$\Delta$-DRESS breaks symmetry by running DRESS on each vertex-deleted subgraph $G\setminus\{v\}$ for every $v\in V$.
+The $\Delta$-DRESS fingerprint is the multiset of per-vertex DRESS fingerprints, or equivalently a pooled histogram accumulating all converged edge values across all $n$ deletions.
+Deleting a vertex from a regular graph produces an irregular subgraph where DRESS can now distinguish structure that was hidden by the uniform regularity.
 6.1 Connection to the Reconstruction Conjecture
-The multiset $\{\!\{\text{DRESS}(G\setminus\{v\}):v\in V\}\!\}$ is directly analogous to the deck in the Kelly–Ulam reconstruction conjecture, which posits that graphs with $n\geq 3$ are determined (up to isomorphism) by their multiset of node-deleted subgraphs.
+The multiset $\{\!\{\text{DRESS}(G\setminus\{v\}):v\in V\}\!\}$ is directly analogous to the deck in the Kelly–Ulam reconstruction conjecture, which posits that graphs with $n\geq 3$ are determined (up to isomorphism) by their multiset of vertex-deleted subgraphs.
 $\Delta$-DRESS computes a continuous relaxation of this deck.
@@
   • Chang Graphs: Distinguished all 6 pairs among T(8) and the three Chang graphs (SRG(28,12,6,4)(28,12,6,4); confound 3-WL).
-  • $2\times C_{4}$ vs. $C_{8}$: Successfully distinguished (both 2-regular on 8 nodes).
-  • Petersen vs. Pentagonal Prism: Successfully distinguished (both 3-regular on 10 nodes).
+  • $2\times C_{4}$ vs. $C_{8}$: Successfully distinguished (both 2-regular on 8 vertices).
+  • Petersen vs. Pentagonal Prism: Successfully distinguished (both 3-regular on 10 vertices).
 7 Family Structure
-The DRESS variants introduced in this paper form a nested hierarchy with an orthogonal composition operator:
+The DRESS variants introduced in this paper form a nested hierarchy with two orthogonal composition operators:
 $\text{Generalized-DRESS}\supset\text{Motif-DRESS}\supset\text{Original-DRESS}$
@@
 $\Delta\text{-DRESS}(\mathcal{F})=\{\!\{\mathcal{F}(G\setminus\{v\}):v\in V\}\!\}$
-The node-deletion strategy is independent of the choice of $\mathcal{N}$, $f$, and $g$.
-All experiments in Section 6 use $\mathcal{F}=\text{Original-DRESS}$, but $\Delta$-Motif-DRESS or $\Delta$-Cosine-DRESS are equally valid and may offer complementary expressiveness.
-The deletion strategy generalizes naturally to $\Delta^{k}$-DRESS, which runs DRESS on all $\binom{n}{k}$ node-deleted subgraphs; we leave the exploration of this hierarchy to future work.
+This strategy is independent of the choice of $\mathcal{N}$, $f$, and $g$.
+$\Delta$-DRESS breaks symmetry by deletion and generalizes to depth $k$: $\Delta^{k}$ runs DRESS on all $\binom{n}{k}$ vertex-deleted subgraphs.
+The exploration of this iterated hierarchy and its relationship to the WL hierarchy is developed in the companion paper [4].
 8 Conclusion
 We presented the DRESS family, a hierarchy of continuous structural refinement algorithms.
-Starting from Original-DRESS—a single equation proposed in 2018—we showed it distinguishes graphs that 1-WL cannot (Theorem 1).
+Starting from Original-DRESS–a single equation proposed in 2018–we showed it distinguishes graphs that 1-WL cannot (Theorem 1).
 We then generalized it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions.
 Generalized-DRESS abstracts further, opening the aggregation function and norm as free parameters.
-$\Delta$-DRESS extends the framework by running DRESS on each node-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture and empirically distinguishing Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs.
+$\Delta$-DRESS extends the framework by running DRESS on each vertex-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture, and empirically distinguishes Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs.
 These results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL, without the prohibitive computational cost of higher-order WL tests.